OpenStudy (anonymous):

What is the solution to the equation? 4x – 6 = –3x + 8 A. B. C. D.

OpenStudy (tkhunny):

A change of attitude will help. 4x – 6 = –3x + 8 Add 6 to both sides 4x – 6 + 6 = –3x + 8 + 6 Simplify 4x = –3x + 14 <== I picked "Add 6" so that it would do that. Add 3x 3x + 4x = 3x – 3x + 14 Simplify 7x = 14 <== I picked "Add 3x" so that it would do that. Divide by 7 7x/7 = 14/7 Simplify x = 2 <== I picked "Divide by 7" so that it would do that. Okay, there is your free one. Now, you do one.